# Relative periodic solutions in spatial Kepler problem with symmetric perturbation

### v1

@@ 4. The ellipsoid problem
In this section, we apply results from the previous section to the ellipsoid problem. Recall that when the eccentricity $\mathfrak{e}$ is samll enough, the potential function can be written as (4). If we further assume $GM=1$, then
@@
    the ellipsoid problem have a solution $\xi_{\varepsilon}(t)$ which locate on $r$-axis and it is linear stable.
@@
Interpreting this system as a Kepler problem with symmetric perturbation, we express the perturbation term as
@@
The existence of $\xi_{\varepsilon}$ follows from Lemma 3.1. Then Theorem 1.3 implies property (b). To establish property (c), we apply Theorem 3.4 through the following calculations.
@@
By Corollary 3.1, $\xi_{\varepsilon}(t)$ is linear stable and this proves property (a). Then we need to verify the condition (ii) in Theorem 3.4.
@@
Then we study the right side,
@@
(57) $\frac{\pi}{4\varpi^{2}}e^{2}(4-3e^{2})<\frac{2\pi C(e)}{\varpi^{2}}(3C(e)-1)\quad\forall e\in(0,1),$
implies condition (c) holds. ∎
@@ 5. The n-pyramidal problem
In this section, we apply the preceding results to the $n$-pyramidal problem. Recall that the fixed axis is $z$-axis and the masses of the $1+n$ bodies are $m_{0},m_{1},\cdots,m_{n}$ and $m_{j}=m$ for $j=1,2,\cdots,n$. Let $q_{1}=(x,y,z)\in\mathbb{R}^{3}$ is the position of $m_{1}$, since the mass center is fixed at the origin, the position of $m_{0}$ is $q_{0}=(0,0,-n\varepsilon z)$, where $\varepsilon=m/m_{0}$ is the mass ratio. Correspondingly, the velocity of $m_{1}$: $\dot{q}_{1}=(\dot{x},\dot{y},\dot{z})$ and $m_{0}$: $\dot{q}_{0}=(0,0,-n\varepsilon\dot{z})$. Then the system can be described by the motion of $m_{1}$ and the Lagrangian is
@@
Through the Legendre transformation
$p_{r}=\dot{r},\quad p_{\theta}=r^{2}\dot{\theta},\quad p_{z}=\dot{z},$
with scaled coordinate $\hat{z}=\sqrt{1+n\varepsilon}\,z$ and fixed angular momentum $r^{2}\dot{\theta}=\varpi$, we obtain the reduced Hamiltonian,
(60) $H_{\varpi,\varepsilon}(p_{r},p_{\hat{z}},r,\hat{z})=\frac{1}{2}(p_{r}^{2}+p_{\hat{z}}^{2})+\frac{\varpi^{2}}{2r^{2}}-\frac{M(n)\varepsilon}{2r}-\frac{1}{\sqrt{r^{2}+(1+n\varepsilon)\hat{z}^{2}}}.$
Hereafter, we denote $\hat{z}$ simply by $z$ when no confusion arises. The parameter set $\Gamma_{\varepsilon}$ can be explicitly calculated for this system.
@@
For $(1+(M(n)\varepsilon)/2)^{2}<2h\varpi^{2}<-(M(n)\varepsilon)^{2}/4$, energy surface $\mathcal{M}_{\varepsilon}$ is a compact manifold homeomorphic to $\mathbb{S}^{3}$ with contact form $\lambda$.
@@
The proof is similar to Proposition 2.1, leaving it to the readers. ∎
@@
    the n-pyramidal problem have a $\varpi$ relative periodic solution $\xi_{\varepsilon}(t)$ which locate on $r$-axis and it is linear stable.
@@
    the n-pyramidal problem admits infinitely many $\varpi$ relative periodic solutions in $\mathcal{M}_{\varepsilon}$.
@@
The linear stability of $\xi_{\varepsilon}(t)$ is refer to the linear stability in this sub-manifold. The global linear stability of this system can see [17],[10],[11] and [20].
@@
We interpret this system as a Kepler problem with symmetric perturbation, where the perturbation term can be written as
@@
where $G(e)=\sqrt{1-\left(\frac{-e}{1+\sqrt{1-e^{2}}}\right)^{4}}$. Here we used the fact that $\int_{0}^{\pi}\frac{\cos nt}{1+e\cos t}dt=\frac{\pi}{\sqrt{1-e^{2}}}\left(\frac{-e}{1+\sqrt{1-e^{2}}}\right)^{n}$ for $e\in(0,1)$. By Corollary 3.1, $\xi_{\varepsilon}(t)$ is linear stable and this implies property (a).
When $2\leq n\leq 472$, this equivalent to $2n>M(n)$, therefor $E(f)>0$ and we need verify condition (ii) in Theorem 3.4. Then we calculate $\tilde{V}(f)$,
@@
Combine above computations, we obtain
@@
Then we calculate the right side,
@@
$2C(e)\tilde{A}(f)+C^{2}(e)\sqrt{D(f)}/2\\ =\frac{\pi}{4}(2n-M(n))\varpi^{2}C^{2}(e)\frac{G(e)}{\sqrt{1-e^{2}}}-\frac{M(n)\pi\varpi^{2}C(e)}{\sqrt{1-e^{2}}}.$
Due to $\frac{G(e)}{\sqrt{1-e^{2}}}>1$ for all $e\in(0,1)$, condition (c) holds. ∎
@@
In this section, we briefly review the index theory for symplectic path. The detail could be found in [15]. Let $(\mathbb{R}^{2n},\omega)$ be the standard symplectic vector space with coordinates $(x_{1},\cdots,x_{n},y_{1},\cdots,y_{n})$, then $\omega=\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i}$. Let $J=\begin{pmatrix}0&-I_{n}\\ I_{n}&0\end{pmatrix}$, where $I_{n}$ is the identity matrix on $\mathbb{R}^{n}$. The symplectic group $\text{Sp}(2n)$ is defined by
@@
Moreover, the $\omega$-regular set of $\text{Sp}(2n)$ is defined by $\text{Sp}(2n)_{\omega}^{*}=\text{Sp}(2n)\setminus\text{Sp}(2n)_{\omega}^{0}$. For any two continuous path $\xi$ and $\eta:[0,\tau]\rightarrow\text{Sp}(2n)$ with $\xi(\tau)=\eta(0)$, we define their concatenation as:

### v2

@@ 4. The ellipsoid problem
In this section, we apply results from the previous section to the ellipsoid problem. Recall that when the eccentricity $\mathfrak{e}$ is small enough, the potential function can be written as (4). If we further assume $GM=1$, then
@@
    the ellipsoid problem has a solution $\xi_{\varepsilon}(t)$ that lies on the $r$-axis and it is linearly stable.
@@
Interpreting this system as a Kepler problem with a symmetric perturbation, we express the perturbation term as
@@
The existence of $\xi_{\varepsilon}$ follows from Lemma 3.1. Theorem 1.3 implies property (b). To establish property (c), we apply Theorem 3.4 through the following calculations.
@@
By Corollary 3.1, $\xi_{\varepsilon}(t)$ is linearly stable and this proves property (a). Then we need to verify condition (ii) in Theorem 3.4.
@@
Then we study the right-hand side,
@@
(57) $\frac{\pi}{4\varpi^{2}}e^{2}(4-3e^{2})<\frac{2\pi C(e)}{\varpi^{2}}(3C(e)-1)\quad\forall e\in(0,1)$
implies condition (ii) holds. ∎
@@ 5. The n-pyramidal problem
In this section, we apply the preceding results to the $n$-pyramidal problem. Recall that the fixed axis is the $z$-axis and the masses of the $1+n$ bodies are $m_{0},m_{1},\cdots,m_{n}$ and $m_{j}=m$ for $j=1,2,\cdots,n$. Let $q_{1}=(x,y,z)\in\mathbb{R}^{3}$ be the position of $m_{1}$, since the mass center is fixed at the origin, the position of $m_{0}$ is $q_{0}=(0,0,-n\varepsilon z)$, where $\varepsilon=m/\penalty 50m_{0}$ is the mass ratio. Correspondingly, the velocity of $m_{1}$ is $\dot{q}_{1}=(\dot{x},\dot{y},\dot{z})$ and that of $m_{0}$ is $\dot{q}_{0}=(0,0,-n\varepsilon\dot{z})$. Then the system can be described by the motion of $m_{1}$ and the Lagrangian is
@@
After a further coordinate change $\hat{z}=\sqrt{1+n\varepsilon}\,z$, the Legendre transformation gives us
$p_{r}=\dot{r},\quad p_{\theta}=r^{2}\dot{\theta},\quad p_{\hat{z}}=\dot{\hat{z}}$,
we obtain the reduced Hamiltonian
(60) $H_{\varpi,\varepsilon}(p_{r},p_{\hat{z}},r,\hat{z})=\frac{1}{2}(p_{r}^{2}+p_{\hat{z}}^{2})+\frac{\varpi^{2}}{2r^{2}}-\frac{M(n)\varepsilon}{2r}-\frac{1}{\sqrt{r^{2}+(1+n\varepsilon)\hat{z}^{2}}}$
with fixed angular momentum $p_{\theta}=r^{2}\dot{\theta}=\varpi$. Hereafter, we denote $\hat{z}$ simply by $z$ when no confusion arises. The parameter set $\Gamma_{\varepsilon}$ can be explicitly calculated for this system.
@@
For $(1+(M(n)\varepsilon)/2)^{2}<2h\varpi^{2}<-(M(n)\varepsilon)^{2}/4$, energy surface $\mathcal{M}_{\varepsilon}$ is a compact contact manifold diffeomorphic to $\mathbb{S}^{3}$ with contact form $\lambda$.
@@
The proof is similar to Proposition 2.1, leaving it to the reader. ∎
@@
    the n-pyramidal problem has a $\varpi$-relative periodic solution $\xi_{\varepsilon}(t)$ that lies on the $r$-axis and it is linearly stable.
@@
    the n-pyramidal problem admits infinitely many $\varpi$-relative periodic solutions in $\mathcal{M}_{\varepsilon}$.
@@
The linear stability of $\xi_{\varepsilon}(t)$ refers to the linear stability in this submanifold. The global linear stability of $\xi_{\varepsilon}(t)$ can be found in [17, 10, 11, 20].
@@
We interpret this system as a Kepler problem with a symmetric perturbation, where the perturbation term can be written as
@@
where $G(e)=\sqrt{1-\left(\frac{-e}{1+\sqrt{1-e^{2}}}\right)^{4}}$. Here we used the fact that $\int_{0}^{\pi}\frac{\cos nt}{1+e\cos t}\mathrm{d}t=\frac{\pi}{\sqrt{1-e^{2}}}\left(\frac{-e}{1+\sqrt{1-e^{2}}}\right)^{n}$ for $e\in(0,1)$. By Corollary 3.1, $\xi_{\varepsilon}(t)$ is linearly stable and this implies property (a).
When $2\leq n\leq 472$, this is equivalent to $2n>M(n)$, therefore $E(f)>0$ and we need verify condition (ii) in Theorem 3.4. Then we calculate $\tilde{V}(f)$,
@@
Combining above computations, we obtain
@@
Then we calculate the right-hand side,
@@
[…]
Due to $\frac{G(e)}{\sqrt{1-e^{2}}}>1$ for all $e\in(0,1)$, condition (ii) holds. ∎
@@
In this section, we briefly review the index theory for symplectic paths. Details can be found in [15]. Let $(\mathbb{R}^{2n},\omega)$ be the standard symplectic vector space with coordinates $(x_{1},\cdots,x_{n},y_{1},\cdots,y_{n})$, then $\omega=\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}y_{i}$. Let $J=\begin{pmatrix}0&-I_{n}\\ I_{n}&0\end{pmatrix}$, where $I_{n}$ is the identity matrix on $\mathbb{R}^{n}$. The symplectic group $\text{Sp}(2n)$ is defined by
@@
Moreover, the $\omega$-regular set of $\text{Sp}(2n)$ is defined by $\text{Sp}(2n)_{\omega}^{*}=\text{Sp}(2n)\setminus\text{Sp}(2n)_{\omega}^{0}$. For any two continuous paths $\xi$ and $\eta:[0,\tau]\rightarrow\text{Sp}(2n)$ with $\xi(\tau)=\eta(0)$, we define their concatenation as:
@@
Data Availability. Data sharing not applicable to this article as no datasets were generated or analyzed during the current study.
Acknowledgement. The authors are deeply grateful to the anonymous referees for their careful reading of the manuscript and insightful comments and suggestions.